What is the value of xwhen f(x) = -1 1) 3 2) -1 3) 5 14 4 o
Answer #1

y i frox) 3 2 > N 3 4 25,-12 A -2 y=fia) -3 say y = frac) fra 2-1 by graph on on point A, a-batue is s so x=5 3 option 3